Factor: x^2 + x^3 + 3x + 3
thank you!
: Factor: x^2 + x^3 + 3x + 3
thank you!

Answer by oscargut(667)

x^2 + x^3 + 3x + 3= x^2(1+x)+3(x+1)=(x^2+3)(x+1)